OpenStudy (anonymous):

Preparing Module 1 (Units and Significant Figures) I don't understand how to get to the step 1.5.7 on page 6: the radius of the standard mass is r = cubic root of V/2 Pi Appreciate anybody's help. Thanks.

OpenStudy (anonymous):

In 1.5.5, r^3 = V/(2 Pi) . Then move the power "3" on left hand side to the right hand sides to become power "1/3" (in other word, cubic root), so r = ( V / (2 Pi) ) ^ (1/3) remark: i use ^ to represent power. e.g. 3^2 = 3 X 3 = 9

OpenStudy (anonymous):

Thank you very much. That was easy. I was confused by the previous step (1.5.6.) which has nothing to do to get to the final result.